Vincent's boat will go 13 miles per hour in still water. If he can go 14 miles downstream in the same amount of time as it takes to go 8 miles upstream, then what is the speed of the current?
Answer by nerdybill(6958)

distance formula:
d = rt
where
d is distance
r is rate or speed
t is time
.
Since we're talking about time, let's rearrange the equation:
d = rt
solving for t:
t = d/r
.
Let x = speed of current
time going downstream
14/(13+x)
time going upstream
8/(13-x)
.
Since "he can go 14 miles downstream in the same amount of time as it takes to go 8 miles upstream"
14/(13+x) = 8/(13-x)
14(13-x) = 8(13+x)
14(13-x) = 8(13+x)
182 - 14x = 104 + 8x
182 - 14x = 104 + 8x
78 = 22x
3.545 mph = x (speed of current)